### v1

# How Effective are Large Time Series Models in Hydrology? A Study on Water Level Forecasting in Everglades

Abstract
The Everglades play a crucial role in flood and drought regulation, water resource planning, and ecosystem management in the surrounding regions. However, traditional physics-based and statistical methods for predicting water levels often face significant challenges, including high computational costs and limited adaptability to diverse or unforeseen conditions. Recent advancements in large time series models have demonstrated the potential to address these limitations, with state-of-the-art deep learning and foundation models achieving remarkable success in time series forecasting across various domains. Despite this progress, their application to critical environmental systems, such as the Everglades, remains underexplored. In this study, we fill the gap by investigating twelve task-specific models and five time series foundation models across six categories for a real-world application focused on water level prediction in the Everglades. Our primary results show that the foundation model Chronos significantly outperforms all other models while the remaining foundation models exhibit relatively poor performance. Moreover, the performance of task-specific models varies with the model architectures. Lastly, we discuss the possible reasons for the varying performance of models.
1 Introduction
The Everglades is a unique ecosystem characterized by vast, swampy grassland, especially containing sawgrass and seasonally covered by slowly moving water. Water levels in the Everglades play a pivotal role in the prevention of floods and droughts, sustainable water resource planning, recreational and infrastructure management and ecosystem management Paudel et al. (2020); Gibble et al. (2020); Saberski et al. (2022). For example, the U.S. Everglades National Park in South Florida, spanning 2,357 square miles, as shown in Figure 1, has a profound influence on urban development, flood control, agricultural expansion, and natural hydrology. Therefore, accurate water level forecasting in the Everglades is of significance, promoting environmental sustainability and benefiting society.
Figure 1: Water Flow through Everglades National Park.
Water levels are typically recorded as time series, and recent advancements in deep learning (DL) have demonstrated its significant potential in time series forecasting. These models excel at capturing nonlinear dependencies across time and variables, enabling both short- and long-term predictions. For example, recurrent neural networks (RNNs) and convolutional neural networks (CNNs) have been used to learn the temporal dependencies for the prediction of water level Shi et al. (2023b). Graph neural networks, often integrated with sequential models, have been utilized to learn complex spatiotemporal relationships from multiple environmental variables Kazadi et al. (2022). More broadly, state-of-the-art models for time series forecasting have demonstrated remarkable performance across diverse domains Liu and Wang (2024), such as electricity, weather, traffic, healthcare, and energy. Most of them leverage Transformer-based architectures Vaswani (2017), which excel at capturing long-term correlations among elements within time series. However, these models usually require high computation costs due to the self-attention scheme with $\mathcal{O}(L^{2})$ time and memory complexity, where $L$ is the sequence length. Although several variants of Transformers Li et al. (2019); Zhou et al. (2021); Nie et al. (2022) have improved the computational complexity to $\mathcal{O}(L\log L)$ or $\mathcal{O}(L)$, significant challenges still persist. Furthermore, some researchers have questioned whether the self-attention mechanism captures temporal information, claiming that even linear models with $\mathcal{O}(L)$ complexity can outperform Transformer-based models for long-term time series forecasting Zeng et al. (2023). Additionally, some researchers have attempted to leverage large language models (LLMs) for time series forecasting Jin et al. (2023). These efforts focus on adapting the comprehensive prior knowledge of LLMs to the time series domain Jin et al. (2024). While these models have demonstrated exceptional performance, they are fully task-specific, requiring retraining from scratch for each new dataset or experimental setting.
Inspired by the success of large language models (LLMs) in natural language processing, researchers have explored the development of foundation models tailored to process time series data. These models adopt a training paradigm similar to that used for LLMs, where each time point in a time series is treated as analogous to a “word” or “token” in a sentence. By leveraging the self-supervised learning technique, these models are pre-trained on large-scale time series datasets to capture complex temporal patterns. The resulting pre-trained models can be fine-tuned to adapt to various downstream tasks. The first time series foundation model is called TimeGPT Garza and Mergenthaler (2023). Other examples include TimesFM Das et al. (2023), Timer Liu et al. (2024a), Moirai Woo et al. (2024) and Chronos Ansari et al. (2024), etc. They have demonstrated remarkable performance in zero-shot setting, reducing or eliminating the need for task-specific retraining.
While these task-specific and foundation models have shown unparalleled performance in time series forecasting across various domains, their application to hydrology is largely unexplored. Here we study the performance of these models for water level prediction in the Everglades, with profound implications to hydrology. Our contributions include:
  • We evaluated 12 task-specific models and 5 time-series foundation models on the accuracy and efficiency of water level prediction in the Everglades.
  • We conducted statistical analyses of prediction errors by assessing models in worst-performing cases, particularly for extremely high and low water levels.
  • We examined the unique features of foundation models, which support varying input lengths and outputs without requiring retraining.
2 Methodology
2.1 Region Selection and Data Collection
The study domain focuses on the Everglades National Park (see Figure 2), including various hydrological structures and stations, and measurements, such as water levels, water flows, precipitation, and evapotranspiration values. We downloaded the corresponding data from the Everglades Depth Estimation Network (EDEN) Haider et al. (2020) and South Florida Water Management District’s Environmental Database (DBHYDRO) South Florida Water Management District (2024). The selection process resulted in a dataset with 37 variables spanning daily records of water levels, water flows, rainfall, and potential evapotranspiration (PET). We provide a summary of the data in Table 3 (see Appendix A).
Figure 2: Study area encompassing portion of the Everglades National Park, covering key hydrological structures and stations.
Figure 3: Flow diagram of the Methodology. a) Data collection: retrieval and merging of raw data. b) Preprocessing: data cleaning, temporal alignment, and interpolation. c) Time-series forecasting (TSF) models (see Section 2.2). d) Evaluation with MAE, RMSE, and SEDI metrics. Details on data collection and pre-processing are provided in Appendix A.
2.2 State-of-The-Art Models
This section briefly introduces task-specific and foundation models for time series forecasting.
2.2.1 Linear-based Models
DLinear and NLinear, introduced in Zeng et al. (2023), are two simple yet effective linear models designed for long-term time series forecasting. DLinear decomposes the raw input data into two components: a trend component, extracted using a moving average kernel, and a remainder (seasonal) component. Each component is processed independently by a single-layer linear model to learn its feature representation, and the two components are then summed up to generate the final prediction. NLinear operates by first subtracting the last value of the input sequence from all elements in the sequence. The adjusted input is then passed through a linear layer to learn the feature representation, after which the subtracted value is added back to generate the final prediction.
2.2.2 MLP-based Models
NBEATS Oreshkin et al. (2019) employs a deep neural architecture based on backward and forward residual links and a deep stack of fully-connected layers. The stacked architecture comprises multiple fully connected blocks, enabling iterative refinement of predictions through backward and forward forecasting. TimeMixer Wang et al. (2024) is based on multi-layer perceptrons (MLPs) and designed for multi-scale time series modeling, capturing both microscopic seasonal and macroscopic trends. TSMixer Chen et al. (2023) is a novel architecture comprising stacked MLPs, which learns from both time and feature dimensions with mixing operations. TSMixerx Chen et al. (2023) is an extension of TSMixer, allowing additional exogenous features, including static features and future information, such as product categories and promotional events.
2.2.3 Transformer-based Models
PatchTST Nie et al. (2022) segments long time series into multiple non-overlapping patches of fixed length, aggregating time steps into subseries-level patches to capture rich local information that is otherwise unavailable at the point level. Additionally, it leverages the Transformer architecture with self-attention mechanisms to model temporal dependencies across these patches, enabling the capture of long-term dependencies while reducing computational overhead. Informer Zhou et al. (2021) addresses the computational challenges of vanilla Transformers in long-horizon time series forecasting by introducing three key innovations. First, it employs a self-attention mechanism, which reduces the time and memory complexity from $\mathcal{O}(L^{2})$ to $\mathcal{O}(L\log L)$ by focusing on the most critical query-key pairs. Second, it incorporates a self-attention distillation process, enabling the model to retain essential temporal dependencies while effectively handling long input sequences. iTransformer Liu et al. (2023) modifies the traditional Transformer architecture by implementing the attention mechanism across multiple variables rather than individual time points, where each time series is embedded as a token. This model has shown promising performance in capturing cross-variable dependencies in multivariate time series data.
2.2.4 KAN-based Models
KANs (Kolmogorov-Arnold Networks) Liu et al. (2024b) are an alternative to MLPs, using the Kolmogorov-Arnold approximation theorem, where splines are learned to approximate more complex functions. Unlike the MLP, the nonlinear functions are learned at the edges. At the nodes, the edge functions are merely summed up. RMoK (Reversible Mixture of KAN) Han et al. (2024b) builds on the KAN concept and employs a mixture-of-experts approach to capture both global and local patterns in time series data. It uses a gating network to assign variables to specialized KAN experts, such as Wav-KAN, JacobiKAN, and TaylorKAN. These experts utilize different mathematical functions and polynomials to effectively model complex time series relationships.
2.2.5 LLM-based Models
TimeLLM Jin et al. (2023) is a reprogramming framework to repurpose LLMs without changes for general time series forecasting. First, it reprograms the input time series with text prototypes by Prompt-as-Prefix (PaP), and then both time series and prompts are fed into a native LLM to align the two modalities. In other words, TimeLLM transforms a forecasting task into a “language task” that can be tackled by an off-the-shelf LLM.
2.2.6 Foundation Models
TimeGPT Garza and Mergenthaler (2023) is the first foundation model pre-trained with enormous amounts of time series data. It is capable of generating accurate zero-shot inference for diverse datasets not seen during training. It is engineered to process time series of different characteristics while accommodating different input sizes and forecasting horizons. TimesFM Das et al. (2023) is a time series foundation model that adopts a decoder-style attention mechanism for pretraining. It processes time series data by segmenting it into patches, enabling TimesFM to capture both local and global temporal dependencies. Timer Liu et al. (2024a) curates large-scale datasets with up to 1 billion time points, unifying heterogeneous time series into a single series into single-series sequence (S3) format, thus developing a GPT-style architecture for large time series models (LTSMs). Moirai Woo et al. (2024) extends the TimesFM architecture to the any-variate setting by “flattening” multivariate time series to a long single sequence. It achieves universal time series forecasting by training a masked encoder-based Transformer to address three challenges inherent in time series data: i) multiple frequencies, ii) multivariate forecasting, and iii) varying distributions. Chronos Ansari et al. (2024) tokenizes time series values using scaling and quantization into a fixed vocabulary and trains the transformer-based architecture via a cross-entropy loss function.
2.3 Forecasting Task
Assuming we are given a dataset where each entry consists of data from $w$ past days (“window size”) including water levels, flows, and other measurements for multiple stations, our goal is to predict water levels for the following $k$ days (“lead time”) at stations of interest. We formulate the forecasting problem as:
$\mathcal{F}_{\theta}(\textbf{X}^{all}_{t-w+1:t})\rightarrow\textbf{Y}^{water}_{t+1:t+k},$ (1)
where $\mathcal{F}_{\theta}$ represents a predictive model with learnable parameters $\theta$. In what follows, we will drop the superscripts and use $\textbf{X}_{t-w+1:t}$ and $\textbf{Y}_{t+1:t+k}$.
Input Features & Output Targets.
We task the model with an instance of supervised regression. To this end, we define the input feature matrix at time step $t$ as $\textbf{X}_{t-w+1:t}=[\textbf{X}_{t-w+1},\dots,\textbf{X}_{t-1},\textbf{X}_{t}]\in\mathcal{R}^{w\times n}$ and the output target as $\textbf{Y}_{t+1:t+k}=[\textbf{Y}_{t+1},\textbf{Y}_{t+2},\dots,\textbf{Y}_{t+k}]\in\mathcal{R}^{k\times m}$, where $n$ and $m$ denote the number of input and output variables. We collected all input-output pairs to form the dataset $\mathcal{D}=\{(\textbf{X}_{t-w+1:t},\textbf{Y}_{t+1:t+k})\}^{T-k}_{t=w}$.
Models Lead Time NP205 P33 G620 NESRS1 NESRS2 Overall
MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE
NLinear Zeng et al. (2023) Linear-based 7 0.195 0.528 0.114 0.508 0.086 0.132 0.081 0.176 0.063 0.113 0.108 0.291
14 0.207 0.313 0.095 0.135 0.108 0.154 0.087 0.134 0.078 0.131 0.115 0.174
21 0.257 0.387 0.120 0.161 0.142 0.189 0.110 0.158 0.102 0.157 0.146 0.210
28 0.323 0.448 0.151 0.189 0.183 0.226 0.138 0.182 0.127 0.179 0.185 0.245
DLinear Zeng et al. (2023) Linear-based 7 0.165 0.245 0.081 0.117 0.094 0.132 0.071 0.118 0.062 0.110 0.095 0.145
14 0.236 0.351 0.125 0.170 0.130 0.174 0.127 0.184 0.127 0.183 0.149 0.212
21 0.352 0.463 0.227 0.297 0.198 0.258 0.266 0.332 0.268 0.331 0.262 0.336
28 0.474 0.598 0.360 0.444 0.278 0.367 0.425 0.500 0.424 0.498 0.392 0.482
NBEATS Oreshkin et al. (2019) MLP-based 7 0.149 0.223 0.062 0.106 0.064 0.116 0.056 0.107 0.049 0.099 0.076 0.130
14 0.195 0.312 0.090 0.140 0.097 0.159 0.083 0.140 0.078 0.138 0.109 0.178
21 0.262 0.415 0.111 0.167 0.123 0.194 0.101 0.167 0.097 0.169 0.139 0.222
28 0.319 0.468 0.139 0.198 0.150 0.227 0.136 0.201 0.133 0.203 0.176 0.259
TimeMixer Wang et al. (2024) MLP-based 7 0.170 0.290 0.079 0.121 0.070 0.126 0.069 0.119 0.061 0.113 0.090 0.154
14 0.249 0.358 0.125 0.168 0.160 0.200 0.109 0.160 0.110 0.163 0.151 0.210
21 0.482 0.692 0.195 0.252 0.248 0.327 0.187 0.239 0.167 0.219 0.256 0.346
28 0.622 0.807 0.228 0.287 0.269 0.350 0.232 0.287 0.208 0.266 0.312 0.399
TSMixer Chen et al. (2023) MLP-based 7 0.145 0.226 0.072 0.111 0.080 0.126 0.063 0.108 0.055 0.100 0.083 0.134
14 0.213 0.328 0.100 0.142 0.115 0.167 0.089 0.135 0.082 0.133 0.204 0.181
21 0.271 0.401 0.126 0.169 0.149 0.202 0.112 0.158 0.106 0.157 0.153 0.217
28 0.339 0.465 0.153 0.200 0.185 0.243 0.130 0.179 0.122 0.173 0.186 0.252
TSMixerx Chen et al. (2023) MLP-based 7 0.366 0.570 0.154 0.205 0.171 0.234 0.139 0.211 0.129 0.204 0.192 0.285
14 0.331 0.405 0.163 0.203 0.205 0.268 0.167 0.230 0.157 0.222 0.204 0.266
21 0.702 1.138 0.242 0.350 0.270 0.398 0.236 0.335 0.222 0.301 0.334 0.504
28 0.560 0.841 0.293 0.409 0.350 0.492 0.312 0.431 0.275 0.368 0.358 0.508
Informer Zhou et al. (2021) Transformer-based 7 0.279 0.389 0.134 0.174 0.160 0.207 0.155 0.193 0.155 0.195 0.177 0.232
14 0.385 0.492 0.262 0.314 0.237 0.311 0.415 0.453 0.423 0.456 0.344 0.405
21 0.438 0.591 0.358 0.403 0.340 0.372 0.152 0.198 0.145 0.195 0.287 0.352
28 0.627 0.781 0.497 0.553 0.658 0.728 0.302 0.363 0.302 0.357 0.478 0.557
PatchTST Nie et al. (2022) Transformer-based 7 0.132 0.208 0.067 0.103 0.071 0.113 0.060 0.105 0.052 0.097 0.076 0.125
14 0.211 0.317 0.096 0.137 0.109 0.156 0.087 0.136 0.079 0.132 0.116 0.176
21 0.280 0.406 0.127 0.168 0.151 0.198 0.114 0.162 0.105 0.161 0.156 0.219
28 0.350 0.480 0.155 0.197 0.191 0.239 0.140 0.188 0.129 0.184 0.193 0.258
iTransformer Liu et al. (2023) Transformer-based 7 0.205 0.320 0.080 0.120 0.083 0.135 0.072 0.118 0.070 0.111 0.102 0.161
14 0.282 0.416 0.121 0.165 0.131 0.184 0.113 0.161 0.105 0.150 0.150 0.215
21 0.310 0.455 0.127 0.175 0.154 0.212 0.129 0.178 0.123 0.171 0.169 0.238
28 0.335 0.503 0.156 0.206 0.193 0.263 0.160 0.215 0.144 0.201 0.198 0.278
KAN Liu et al. (2024b) KAN-based Model 7 0.176 0.268 0.085 0.131 0.101 0.146 0.076 0.130 0.071 0.123 0.102 0.160
14 0.257 0.374 0.126 0.169 0.148 0.203 0.116 0.168 0.110 0.165 0.151 0.216
21 0.292 0.441 0.144 0.196 0.160 0.225 0.138 0.200 0.133 0.199 0.174 0.252
28 0.357 0.502 0.165 0.216 0.188 0.258 0.185 0.243 0.178 0.240 0.214 0.292
RMoK Han et al. (2024b) KAN-based Model 7 0.171 0.247 0.088 0.118 0.100 0.137 0.080 0.116 0.070 0.106 0.102 0.145
14 0.264 0.365 0.129 0.160 0.152 0.190 0.110 0.148 0.099 0.139 0.151 0.200
21 0.310 0.438 0.145 0.184 0.179 0.226 0.133 0.172 0.124 0.170 0.178 0.238
28 0.337 0.483 0.153 0.202 0.185 0.250 0.143 0.194 0.134 0.185 0.191 0.263
TimeLLM Jin et al. (2023) LLM-based Model 7 0.228 0.308 0.106 0.135 0.125 0.160 0.090 0.124 0.080 0.121 0.126 0.170
14 0.311 0.440 0.143 0.177 0.185 0.221 0.130 0.171 0.118 0.169 0.178 0.236
21 0.349 0.489 0.166 0.200 0.211 0.248 0.146 0.189 0.132 0.184 0.201 0.262
28 0.413 0.542 0.196 0.232 0.258 0.298 0.179 0.217 0.162 0.206 0.242 0.299
TimeGPT Garza and Mergenthaler (2023) Foundational Model 7 0.160 0.262 0.082 0.138 0.086 0.153 0.074 0.136 0.065 0.131 0.093 0.164
14 0.247 0.400 0.111 0.165 0.120 0.181 0.102 0.167 0.090 0.165 0.134 0.216
21 0.298 0.492 0.143 0.217 0.159 0.246 0.132 0.215 0.126 0.222 0.172 0.278
28 0.410 0.608 0.204 0.263 0.229 0.310 0.179 0.259 0.166 0.273 0.238 0.342
TimesFM Das et al. (2023) Foundational Model 7 0.224 0.337 0.118 0.166 0.121 0.186 0.108 0.171 0.094 0.156 0.133 0.203
14 0.376 0.584 0.182 0.244 0.177 0.269 0.153 0.229 0.154 0.230 0.208 0.311
21 0.536 0.759 0.246 0.310 0.256 0.342 0.211 0.282 0.206 0.282 0.291 0.395
28 0.639 0.869 0.291 0.350 0.299 0.381 0.246 0.316 0.234 0.301 0.342 0.443
Timer Liu et al. (2024a) Foundational Model 7 0.293 0.374 0.160 0.194 0.181 0.224 0.146 0.184 0.131 0.171 0.182 0.229
14 0.456 0.594 0.238 0.277 0.291 0.344 0.212 0.251 0.193 0.235 0.278 0.340
21 0.559 0.720 0.293 0.330 0.386 0.433 0.265 0.299 0.230 0.270 0.347 0.410
28 0.611 0.784 0.322 0.366 0.446 0.501 0.294 0.331 0.251 0.291 0.385 0.455
Moirai Woo et al. (2024) Foundational Model 7 0.471 0.742 0.205 0.299 0.285 0.424 0.186 0.184 0.161 0.231 0.261 0.389
14 0.483 0.755 0.212 0.305 0.294 0.429 0.187 0.264 0.165 0.242 0.265 0.397
21 0.551 0.817 0.328 0.418 0.432 0.557 0.255 0.343 0.223 0.309 0.358 0.489
28 0.555 0.820 0.331 0.422 0.446 0.569 0.261 0.351 0.224 0.305 0.364 0.493
Chronos Ansari et al. (2024) Foundational Model 7 0.094 0.132 0.047 0.059 0.042 0.054 0.036 0.051 0.028 0.045 0.049 0.075
14 0.127 0.176 0.069 0.089 0.060 0.073 0.053 0.073 0.037 0.058 0.069 0.103
21 0.148 0.198 0.087 0.107 0.074 0.089 0.069 0.089 0.047 0.066 0.085 0.119
28 0.147 0.200 0.090 0.116 0.083 0.097 0.071 0.095 0.049 0.068 0.088 0.124
Table 1: Model performance for 5 stations (NP205, P33, G620, NESRS1, and NESRS2) for lead times of 7, 14, 21, and 28 days. The first 12 models are task-specific while the last 5 are pre-trained foundation models for time series modeling. The best results are in bold.
3 Experiments
3.1 Experimental Setup
Dataset splitting.
The data was split into training, validation, and test sets with an approximate ratio of $70\%:15\%:15\%$. The training set includes data from October 16, 2020, to July 1, 2023, representing 989 days. The validation set spans July 2, 2023, to January 28, 2024, covering 211 days, and the test set includes data from January 29, 2024, to August 26, 2024, also comprising 211 days.
Task Settings.
To ensure consistent evaluation across different models, we considered all variables from the previous 100 days as input and trained all task-specific models to forecast water levels at 5 stations (NP205, P33, G620, NESRS1, and NESRS2, as shown in Figure 2) for the next 7, 14, 21, and 28 days. Time Series Foundation Models are with the same setting in zero-shot inference (without training).
Evaluation Metrics.
We used Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) to evaluate the overall performance of these models. MAE measures the predictive robustness of an algorithm but is relatively insensitive to outliers, whereas RMSE can amplify errors. Additionally, we also include the Symmetric Extremal Dependence Index (SEDI) as a specialized metric for extreme values. It classifies each prediction as either “extreme” or “normal” based on upper or lower quantile thresholds. SEDI values belong to the range $[0,1]$, where higher values indicate better accuracy in identifying extreme water levels.
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}\left|y_{i}-\hat{y}_{i}\right|,\text{and}$ (2)
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}-\hat{y}_{i}\right)^{2}},\text{and}$ (3)
$\text{SEDI}(p)=\frac{\Sigma(\hat{y}<y_{\text{low}}^{p}\&y<y_{\text{low}}^{p})+\Sigma(\hat{y}>y_{\text{up}}^{p}\&y>y_{\text{up}}^{p})}{\Sigma(y<y_{\text{low}}^{p})+\Sigma(y>y_{\text{up}}^{p})},$ (4)
where $y$ and $\hat{y}$ denote the ground truth and prediction, respectively. The tests, $\hat{y}<y_{\text{up}}^{p}$ and $y<y_{\text{up}}^{p}$ judge whether the predicted and observed values are extremes based on the threshold value $y_{\text{up}}^{p}$. Following the work Han et al. (2024a), we set $10\%$ and $90\%$ as the low and high thresholds, respectively. $N$ is the number of samples.
Implementation Details.
For task-specific models, we first pre-train them before conducting inference. The training is performed for a total of 1,000 epochs, starting with a learning rate of $1e^{-3}$. We used a batch size of 32 for training, and early stopping is executed if training loss does not decrease for 50 iterations. For pre-trained time series foundation models, we directly used their existing model weights for zero-shot inference on our test set. Experiments were conducted on an NVIDIA A100 GPU with 80GB memory. More details are provided in Appendix F.
4 Results
Table 1 compares the quantitative performance of 12 task-specific and 5 foundation models across 5 stations and over 4 lead times, allowing us to address the following questions.
RQ1: How do time series foundation models perform compared to task-specific models?
Surprisingly, the time series foundation model, Chronos, consistently outperforms other models over various lead times across all stations, showcasing exceptional performance in zero-shot inference. Temporally, while the performance of all models declines as the lead time increases, Chronos demonstrates a comparatively slower drop. Spatially, we observe the performance of all models varies across stations with the worst scenarios for the NP205 station, where Chronos still achieves the lowest prediction errors. A possible reason for the lower performance at the NP205 station is its weaker correlation with other water stations (see Figure 8 in Appendix D). Note that the remaining foundation models show relatively poor performance. TimeGPT achieves results comparable to the best task-specific models, NBEATS and PatchTST, while TimesFM, Timer and Moirai perform much worse. Appendix E includes the qualitative analysis to provide a clear observation.
RQ2: How do the task-specific models perform?
The performance varies across their model architectures, with NBEATS, TSMixer, PatchTST, and RMoK consistently outperforming others within their respective categories. RMoK model outperforms the vanilla KAN, particularly for long-term forecasting, highlighting the effectiveness of the mixture-of-experts approach. Furthermore, the linear-based models excel in short-term predictions but experience a significant drop in performance as the forecasting horizon increases. In contrast, MLP-based NBEATS and Transformer-based PatchTST help mitigate this issue.
RQ3: How do the 17 models perform with respect to extreme values?
Following the work for extreme weather prediction Han et al. (2024a), we use the SEDI metric to evaluate model performance under extreme values and report the results in Table 2. We observed that Chronos significantly outperforms other models. Moreover, even though TSMixerx shows average performance in overall water level prediction (see Table 1), it performs notably well for predicting extreme water levels at the P33, G620, and NESRS1 stations. To better understand this observation, we visualize the ground truth and model predictions with a lead time of 28 days (see Figure 6 in Appendix B). The visualizations reveal that TSMixerx generally follows the global trend but fails to capture the temporal details (e.g., abrupt changes around time points 50, 130, 150). Notably, despite producing several false alarms between time 140 and 170, TSMixerx successfully predicts most extreme values highlighted in red circles.
Models NP205 P33 G620 NESRS1 NESRS2 Overall
NLinear 0.143 0.122 0.146 0.122 0.167 0.140
DLinear 0.262 0.512 0.390 0.488 0.500 0.430
NBEATS 0.262 0.268 0.463 0.268 0.286 0.310
TimeMixer 0.119 0.049 0.341 0.000 0.000 0.102
TSMixer 0.143 0.146 0.195 0.073 0.095 0.131
TSMixerx 0.381 0.829 0.829 0.780 0.476 0.659
Informer 0.500 0.488 0.487 0.512 0.500 0.498
PatchTST 0.190 0.146 0.195 0.146 0.190 0.174
iTransformer 0.143 0.244 0.244 0.244 0.595 0.294
KAN 0.119 0.073 0.341 0.171 0.190 0.179
RMoK 0.119 0.098 0.171 0.317 0.309 0.203
TimeLLM 0.001 0.001 0.001 0.001 0.001 0.001
TimeGPT 0.024 0.195 0.195 0.171 0.167 0.150
TimesFM 0.095 0.024 0.170 0.000 0.000 0.058
Timer 0.000 0.000 0.000 0.000 0.000 0.000
Moirai 0.476 0.317 0.488 0.122 0.023 0.285
Chronos 0.786 0.659 0.805 0.537 0.762 0.710
Table 2: SEDI values (the higher, the better) for predicting extreme values across 5 stations with lead time of 28 days.
RQ4: How do models perform with different model sizes?
Figure 4 compares models by visualizing their prediction accuracy (MAE), inference time, and model complexity. We found that while the model size influences predictive accuracy and inference time, it does not solely determine overall performance. For task-specific models, despite having a smaller size, NBEATS outperforms Informer and TimesLLM, both of which are larger. A similar phenomenon is observed between Chronos and other foundation models. However, when comparing task-specific models to foundation models, we found that only Chronos significantly surpasses task-specific models in performance. This suggests that simply scaling model parameters may not improve prediction accuracy, guiding researchers to explore alternative strategies to optimize foundation models, such as expanding datasets covering more domains and designing more efficient network architectures.
Figure 4: Plot of Accuracy vs Efficiency vs Model Size for an input length of 100 days and lead time of 28 days.
RQ5: What makes Chronos the highest performing model?
While one could attribute the superior performance of Chronos to its tokenization via scaling and quantization or the use of cross entropy in training, a simpler explanation could be that it was exposed to training data that correlates with data from the Everglades National Park. Even though there are no actual Everglades data included, we found that multiple weather datasets were used for its pre-training Ansari et al. (2024). One of them contains daily measurements of precipitation, snow, snow depth, and temperature from climate stations located in 48 states in the USA. However, the other four foundation models in this study were not pre-trained on this dataset or similar datasets.
RQ6: Does input length impact the predictive accuracy of the models?
One key advantage of foundation models is their ability to generate predictions for inputs of varying lengths without being retrained. This differs from task-specific models, which require retraining once the input lengths change. To analyze the impact of input length on prediction performance, we performed experiments predicting 28 days ahead with varying input lengths and reported the results in Figure 5. It shows that the prediction accuracy depends on the input lengths. The MAE values drop as the input length increases from 25 to 100 days, after which the performance stabilizes.
Figure 5: Performance (lead time: 28 days) vs Input Length. We compute the MAE values of ten samples for each input length.
5 Related Work
Physics-based Models.
Physics-based methods have long been a cornerstone in hydrologic and hydraulics research. For example, 1D and 2D HEC-RAS models have been used to predict water levels Muñoz et al. (2022) and to simulate sewer and street flows in urban systems Huang and Lee (2023). The South Florida Water Management Model (SFWMM) serves as a regional hydrological model designed to simulate water levels in the Everglades SFWMD (2005). Furthermore, researchers have also combined physics-based approaches with ML techniques, such as integrating a hydrologic-hydrodynamic coupling (H2C) model with long short-term memory (LSTM) networks to achieve highly accurate water level predictions in tidal reaches Fei et al. (2023). Despite their utility, these mechanistic models require complex variable integration and rely on predefined assumptions.
Statistical Models.
The Everglades Forecasting application (EverForecast) Pearlstine et al. (2020) is a statistical model designed to generate spatially continuous simulations of the Everglades water levels in South Florida. Autoregressive Integrated Moving Average (ARIMA) model has been used for water level and discharge predictions Agaj et al. (2024), while the Seasonal Autoregressive Integrated Moving Average (SARIMA) model Azad et al. (2022) was explored for forecasting Malaysia’s mean sea level at coastal stations. In Sekban et al. (2022), the study focuses on predicting monthly dam water levels by comparing various statistical time series models. These models excel at identifying patterns from historical data but are limited in capturing complex nonlinear interactions and dependencies among variables.
Data-driven Machine Learning Models.
ML models such as support vector machines (SVM) Khan and Coulibaly (2006), and extreme gradient boosting (XGBoost) Nguyen et al. (2021) have demonstrated the ability to model nonlinear relationships in water level datasets. We have also experimented with these two ML models, but the results were relatively poor. Besides, the advent of deep learning has revolutionized hydrological forecasting, learning complex spatiotemporal patterns from large datasets. Artificial neural networks (ANNs) Abiy et al. (2022), convolutional neural networks (CNNs) Bassah et al. (2025), recurrent neural networks (RNNs) Shi et al. (2023b), Transformer models Hou et al. (2024), and many hybrid models Shi et al. (2023a) have been used to capture temporal dependencies in hydrological data. Furthermore, Transformer-based models, such as Informer Zhou et al. (2021) and PatchTST Nie et al. (2022), are proposed for time series modeling across various domains. NLinear and DLinear are two models with only linear layers Zeng et al. (2023). Neural Basis Expansion Analysis for Interpretable Time Series Forecasting (NBEATS) Oreshkin et al. (2019) and TimeMixer Chen et al. (2023) are two MLP-based architectures for multi-scale time series modeling. Leveraging large language models (LLMs) for time series forecasting was investigated in Jin et al. (2024), where a time series was treated as a sequence of tokens (values) and pre-trained LLMs were fine-tuned with well-designed prompts Jin et al. (2023).
Time Series Foundation Models.
TimeGPT Garza and Mergenthaler (2023), a pioneering foundation model, employs a transformer-based architecture and was pre-trained on massive time series datasets, supporting both fine-tuning and zero-shot inference. TimesFM Das et al. (2023) follows a similar approach, utilizing a decoder-based architecture to improve time series forecasting. Moment Goswami et al. (2024) introduces a multi-scale learning mechanism, enhancing its ability to model heterogeneous temporal dynamics. Timer Liu et al. (2024a) leverages a hybrid structure incorporating state-space models to improve long-term forecasting stability. Moirai Woo et al. (2024) unifies multiple time series forecasting paradigms within a single foundation model framework. Finally, Chronos Ansari et al. (2024) introduces an innovative design that is optimized for zero-shot forecasting across numerous time series applications.
6 Discussion
We evaluated 12 task-specific deep learning models and 5 pre-trained time series foundation models for a practical hydrological application, i.e., water level prediction in the Everglades. For task-specific models, performance varies based on architecture and training. They require careful retraining whenever datasets or experimental settings change. For those foundation models, not all of them demonstrate strong performance. Chronos with the zero-shot inference achieves the best performance while others perform poorly. In addition to model architectures, a potential explanation is that Chronos has included training data that highly correlates with data from the Everglades while the training data sets for other foundation models have a low correlation with data from the Everglades or provide limited representations of the hydrological applications. Therefore, incorporating comprehensive data sets for pre-training is crucial to enhance the performance of foundation models. Furthermore, all models struggle to accurately predict extreme high and low values especially when abrupt changes appear. This limitation suggests that the models may have difficulty capturing sudden shifts in the underlying dynamics, potentially due to insufficient training samples for extreme events or inherent model biases. Future improvements could involve incorporating additional physical constraints Yin et al. (2023), leveraging ensemble methods Wu et al. (2023), or integrating uncertainty quantification Shi et al. (2024) to enhance robustness in extreme conditions. Lastly, model performance varies across various locations, demanding context-aware, region-specific adaptions to improve prediction performance Li et al. (2024). These observations further highlight the critical role of domain knowledge in establishing essential physical constraints for more reliable and accurate predictions.
7 Conclusions
In conclusion, our findings highlight the potential applications of time series foundation models in hydrology. The strong performance of Chronos suggests that certain pre-trained models can leverage their learned representations effectively, even in specialized applications. Moreover, the flexibility of zero-shot inference across different input lengths underscores the high adaptability of foundation models, making them attractive for real-world deployment where data availability and sequence lengths vary. However, further studies are needed to fully harness the potential of these foundation models for hydrologic applications. Exploring strategies to bridge the gap between general-purpose pre-training and the specific requirements of environmental forecasting presents a promising direction for future research.
Collaboration and Broader Impact
ML and Everglades researchers were involved in the multidisciplinary collaborations. The domain experts helped in area selection, data acquisition, data preprocessing, and analysis of the results. This work helps domain researchers select the best AI models for water level forecasting in the Everglades and provides practical insights for government agencies interested in using AI-driven hydrological applications.
Appendix
Appendix A Data Set
A.1 Data summary
Table 3 provides a summary of the key features included in the dataset used for the study. The data encompasses multiple hydrological and meteorological variables collected across various measuring stations. Daily rainfall data, measured in inches, is available for two measuring stations, namely NP205 and P33. Potential Evapotranspiration (PET), measured in millimeters, represents the amount of water that would evaporate and transpire under normal conditions. PET is critical for understanding the water balance in the region as it helps to estimate the loss of water due to evaporation and transpiration. Daily pump and gate flow data is measured in cubic feet per second (cfs). Pumps and gates are used to regulate the water flow between different parts of the Everglades system, making the gate flow data essential for modeling water management operations. Daily water level data, recorded in feet is available for fourteen stations.
Feature Interval Unit #Var. Location
Rainfall Daily $inches$ 2 NP-205, P33
Potential Evapotranspiration Daily $mm$ 2 NP-205, P33
Pump flows Daily $cfs$ 8 S356, S332B, S332BN, S332C, S332DX1, S332D, S200, S199
Gate flows Daily $cfs$ 11 S343A, S343B, S12A, S12B, S12C, S12D, S [^3] N, S333, S334, S344, S18C
Water levels Daily $ft$ 14 S12A, S12B, S12C, S12D, S333, S334, NESRS1, NESRS2, NP-205, P33, G620, EVER4, TSH, NP62
Table 3: Summary of the dataset.
A.2 Pre-processing
Data Cleaning. The raw water level data was reported in two different vertical datums: the North American Vertical Datum of 1988 (NAVD88) and the National Geodetic Vertical Datum of 1929 (NGVD29). Since direct comparisons require a consistent datum, adjustments were made following domain expert recommendations to standardize all values to NGVD29. For example, the G620 station requires an adjustment of +1.51 ft, while S [^3] T requires +1.54 ft, among others. These values were carefully selected based on the suggestions from domain experts. Temporal Alignment. The data pre-processing ensured temporal consistency across all stations. Daily data points were synchronized to a uniform timestamp. Data Interpolation. Completeness of data is vital for accurate forecasting of water levels, particularly in ecologically sensitive regions such as the Everglades National Park (ENP). There are certain gates and stations that still exhibit missing values. We employed time-based interpolation and backward fill techniques to address the missing data.
Appendix B Extreme Value Predictions
Figure 6 highlights the extreme value predictions for the Chronos and TSMixerx models across the P33, G620, and NESRS1 measuring stations for a lead time of 28 days. The two horizontal lines in each plot represent the upper and lower thresholds for extreme values. Predictions that exceed these thresholds are circled, indicating instances where the models detected extreme water levels. Chronos consistently outperforms other models in extreme value prediction, while TSMixerx, despite its ordinary performance for overall water levels, successfully captures extreme events at these three stations. However, we observe false alarms in certain time periods, particularly between time steps 140 and 170. During this window, TSMixerx incorrectly predicts extreme water levels, which could potentially lead to unnecessary warnings or misinterpretations of water conditions.
Figure 6: Extreme value predictions at (a) P33, (b) G620, and (c) NESRS1 stations using Chronos and TSMixerx models.
Appendix C Parameter Study for Input Length
As shown in Figure 7, we investigated how the Chronos model’s predictive performance is affected by varying the input sequence length. Specifically, we tested input lengths of 25, 50, 75, 100, 125, and 150 days, and measured the resulting Mean Absolute Error (MAE) across five different measuring stations (NP205, P33, G620, NESRS1, and NESRS2). The results show that as the number of input days increases, the model’s performance improves, as evidenced by the decreasing MAE values. This indicates that longer input sequences help the model make more accurate predictions. Interestingly, the NP205 measuring station exhibited slightly higher MAE values compared to the other stations, indicating a potential anomaly or unique characteristic of this station. On the other hand, the other stations (P33, G620, NESRS1, and NESRS2) followed a similar trend, with their MAEs closely clustered and showing a consistent improvement as the input sequence length grew.
Figure 7: Impact of Input Length on Chronos Model Performance at various stations.
Appendix D Correlation Study
To investigate the discrepancy between the NP205 station and other water stations, we compute their correlations. The right plot emphasizes correlations among NP205, P33, G620, NESRS1, and NESRS2, while the left plot extends this analysis to include nearby gates. The results reveal that NP205 exhibits lower correlations with the other stations, with values ranging from 0.76 to 0.85, whereas the remaining stations maintain higher correlations (at least 0.85). This aligns with the spatial distribution shown in Figure 2, where NP205 is positioned separately from the others. The hydrological flow dynamics, influenced by water management structures and natural flow patterns from WCA-3A and WCA-3B to Florida Bay, further explain these variations.
Figure 8: Correlation Analysis of water levels at five measuring water stations and nearby gates.
Appendix E Visualization
In the following, we provide the visualizations of the ground truth and prediction values of 16 models at different stations, while the corresponding MAE values are included in the figures.
E.1 Details for the NP205 Station
The model comparison plots shown in Figure 9 visualize the 28-day lead time water level forecasts at NP205, comparing actual observations with predictions from multiple machine learning models. Each subplot highlights predicted trends alongside actual values while also displaying the Mean Absolute Error (MAE), a key metric for evaluating predictive accuracy. The results indicate that Chronos consistently outperforms other models, achieving the lowest MAE of 0.147. This is consistent with the results shown in Table 1, which comprehensively evaluates the performance of the models at multiple measuring stations. The analysis underscores the effectiveness of specialized time series models and pretrained foundation models in capturing hydrological patterns for improved water level forecasting.
Figure 9: Comparison of Model Performance for 28-Day Water Level Forecasting at NP205 Measuring Station.
E.2 Details for the P33 Station
Next, the model comparison plots shown in Figure 10 visualize the 28-day lead time water level forecasts at P33 Measuring Station, comparing actual observations with predictions from multiple models. Unlike NP205, where the models struggled with higher MAE values, all models performed better for the P33 measuring station, achieving lower MAE values across the board. The Chronos model again emerged as the most accurate, achieving the lowest MAE of 0.090, reinforcing its superior ability to capture hydrological patterns. In particular, the NBEATS model, which had an MAE of 0.319 for NP205, improved significantly at P33, achieving a much lower MAE of 0.139.
Figure 10: Comparison of Model Performance for 28-Day Water Level Forecasting at P33 Measuring Station.
E.3 Details for the G620 Station
Figure 11 presents the 28-day lead time water level forecasts for G620, comparing the actual observations with predictions from multiple models. From the analysis, Chronos and DLinear performed notably better compared to their metrics for NP205 and P33. Chronos achieved the lowest MAE of 0.083, while DLinear improved significantly to 0.278 from its previous 0.360 at P33. However, other models exhibited slightly higher MAE values compared to P33. In particular, the Informer model showed a decline in performance, with its MAE increasing to 0.658 from 0.497 at P33, indicating a greater deviation of predicted values from actual observations.
Figure 11: Comparison of Model Performance for 28-Day Water Level Forecasting at G620 Measuring Station.
E.4 Details for the NESRS1 Station
Figure 12 illustrates the 28-day lead-time water level forecasts at the NESRS1 measuring station, comparing actual observations with model predictions. In all models, performance improved significantly compared to the measuring stations NP205, P33, and G620, with lower MAE values indicating improved predictive accuracy. However, an exception is observed with the DLinear model, which recorded an MAE of 0.425, higher than the 0.278 and 0.360 values at the G620 and P33 stations respectively. This suggests that while most models effectively capture water level patterns for NESRS1, DLinear struggles to maintain its prior performance. The Chronos model achieved the best performance with an MAE of 0.071, the lowest among NP205, P33, and G620 measuring stations, highlighting its effectiveness in capturing water level patterns at NESRS1.
Figure 12: Comparison of Model Performance for 28-Day Water Level Forecasting at NESRS1 Measuring Station.
E.5 Details for the NESRS2 Station
Figure 13 presents the 28-day lead time water level forecasts at NESRS2, demonstrating a consistent improvement in predictive accuracy across all models compared to the other measuring stations. The results indicate that NESRS2 exhibits the lowest overall MAE values, making it the most predictable among the five stations analyzed. As observed in prior analyses, the Chronos model continues to outperform others, achieving an MAE of 0.049, the lowest recorded across all stations. Additionally, TSMixerx improved significantly, reducing its MAE from 0.312 at NESRS1 to 0.275 at NESRS2. Similarly, NLinear also demonstrated a slight improvement, with its MAE decreasing from 0.138 to 0.127. These results highlight the effectiveness of specialized time series models in capturing hydrological patterns at NESRS2, further reinforcing the trend of improved predictions at this location.
Figure 13: Comparison of Model Performance for 28-Day Water Level Forecasting at NESRS2 Measuring Station.
Appendix F Implementation Details
The implementation of the forecasting models in this study spans a variety of architectures, ranging from linear models to deep learning-based transformers,large language models and foundation models. Table 4 provides a comprehensive comparison of key implementation parameters, including input size, learning rate, optimizer, library/API dependencies, architecture, training and inference strategies. Additionally, model complexity in terms of the number of parameters is depicted in Figure 4.
Model Input Size Learning Rate Optimizer Library/API Architecture Training Inference
NLinear 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast Linear Yes zero-shot
DLinear 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast Linear Yes zero-shot
NBEATS 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast MLP Yes zero-shot
TimeMixer 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast MLP Yes zero-shot
TSMixer 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast MLP Yes zero-shot
TSMixerx 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast MLP Yes zero-shot
Informer 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast Transformer Yes zero-shot
PatchTST 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast encoder-only Yes zero-shot
iTransformer 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast encoder-only Yes zero-shot
KAN 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast KAN Yes zero-shot
RMok 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast KAN Yes zero-shot
TimeLLM 100 $1e^{-3}$ Adam NeuralForecast [^1] https://github.com/Nixtla/neuralforecast LLM Yes zero-shot
TimeGPT 100 $1e^{-4}$ Adam Nixtla API [^2] https://docs.nixtla.io/docs/getting-started-about_timegpt Transformer No zero-shot
TimesFM 128 $1e^{-4}$ Adam HuggingFace [^3] https://huggingface.co/google/timesfm-1.0-200m-pytorch decoder-only No zero-shot
Timer 100 $1e^{-4}$ Adam HuggingFace [^4] https://huggingface.co/thuml/timer-base-84m decoder-only No zero-shot
Moirai 100 $1e^{-3}$ Adam HuggingFace [^5] https://huggingface.co/Salesforce/moirai-1.0-R-large encoder-only No zero-shot
Chronos 100 $1e^{-5}$ Adam HuggingFace [^6] https://huggingface.co/autogluon/chronos-bolt-small encoder-decoder No zero-shot
Table 4: Implementation Details. The first 12 models are task-specific while the last 5 are pre-trained foundation models for time series modeling.
The NeuralForecast library, part of the Nixtlaverse ecosystem, provides a scalable and user-friendly framework for implementing deep learning models for time series forecasting. In this study, several models from this library were utilized, including NLinear, DLinear, NBEATS, TimeMixer, TSMixer, TSMixerx, Informer,PatchTST, iTransformer, KAN, RMok and TimeLLM, as listed in Table 4. NeuralForecast requires input data in a long-format structure, where each time series is stored as individual rows in a Pandas DataFrame, with corresponding timestamps and values. This format ensures compatibility across multiple time series and facilitates batch processing for neural network-based models. TimeGPT, which is Nixtla’s generative pre-trained transformer for time series forecasting, was evaluated using the Nixtla API, enabling seamless model inference without requiring fine-tuning.
While most models in Table 4 use an input size of 100, TimesFM is assigned an input length of 128, deviating from the standard. The architectural constraint of TimesFM requires input lengths to be multiples of 32. For this study, the TimesFM-1.0-200M checkpoint (timesfm-1.0-200m-pytorch) was used, which supports univariate time series forecasting with context length up to 512. The Timer model was implemented using the Hugging Face transformers library, leveraging its pre-trained Timer-base-84M checkpoint. This model follows a Causal Transformer (Decoder-only) architecture and is loaded using the AutoModelForCausalLM class, enabling generative forecasting for time-series data. Salesforce’s Uni2TS, a PyTorch-based library was used to evaluate the Moirai model for time series forecasting. Inference was conducted in a zero-shot setting without fine-tuning, leveraging the GluonTS framework. The Chronos model was implemented using the AutoGluon-TimeSeries (AG-TS) library, which provides an easy-to-use framework for time series forecasting. The model was accessed through the TimeSeriesPredictor API, specifying the ”amazon/chronos-bolt-small” checkpoint for inference.